# On the diversity of strongly-interacting Type IIn supernovae

### v1

@@ 2.6 Analysis on the line profile and emission
During the interaction between the SN ejecta and the CSM, four main region should be considered: from the outside, i) the unshocked CSM, ii) the CSM that was shocked by a forward shock (FS), iii) the SN ejecta shocked by the reverse shock (RS), and iv) the unshocked SN ejecta that are expanding fast (Chevalier & Fransson, 1994). If the CSM is dense (), a cool dense shell (CDS) forms between the FS and RS (Chevalier & Fransson, 1985). An opaque CDS has the same effect of an expanding photosphere (Chugai, 2001). In the case of strong interaction, the heated CSM dominates the optical emission. Since emissions from the CSM can remain bright for months or years, it is possible that the ejecta thermal energy fades before the CSM becomes transparent and thus one may never see the characteristic P-Cygni profiles indicative of an expanding photosphere (Smith, 2017). In general, in SNe IIn one may expect a narrow () component due to the unshocked CSM and a broader () component which, depending on the optical depth of the CSM can originate either from the ionised SN ejecta or from electron-scattering in the CSM (Huang & Chevalier, 2018). Furthermore, if the CSM is asymmetric (Stritzinger et al., 2012), or if dust is present (Fox et al., 2011), they can affect the shape of the line profiles, thus making the theoretical interpretation more difficult.
@@
In Fig. 11, we plot a zoom on the H region of spectra taken around 100 from the explosion for all the SNe in our sample. SNe 2021adxl and 2022qml have broad, blue, asymmetric profiles with a narrow emission centered at rest wavelength. On the other hand, SNe 2021acya and 2022wed have a more symmetric broad line (see Tabs. 14–21). In the case of SN 2021acya, there is a small dip that may be a narrow P-Cygni absorption but it is probably an instrumental artifact, since it is not visible at other epochs.
@@
Given the diversity in shape of the H across our sample, it is difficult to directly compare them in terms of FWHM, line position, and total flux. Opting to apply the same method to all of them while being conscious that the procedure will not be optimal, we performed a multi-component fit for the H lines, including the contribution of a broad Gaussian and a narrow Lorentzian function. We chose these rather simple profiles because they can ensure a more direct comparison among all the SNe in the sample, even if they might not provide the most perfect fit, especially for complicated profiles such as those of SNe 2021adxl and 2022qml. In some spectra, adding a second, intermediate-width Gaussian component allowed for a better fit. This is a relatively faint feature that was ignored for the rest of the analysis. For each component, we fit the position, peak intensity, and width. Given the asymmetric profile of the line in some objects, the position of the peak of the broad component was allowed to vary with respect to the narrow one. All the measurements were performed on reddening-corrected spectra.
@@
The measured fit parameters as a function of the phase from the explosion are reported in Tabs. 14–21. For all the SNe in the sample, position, FWHM, and intensity of the narrow component are constant within the errors (calculated by summing in quadrature the RMS and the resolution). Also, the FWHM of the narrow component is not resolved in the fit with the exception of the spectrum of SN 2022qml at +49, which yields a FWHM of 12.9 Å. The average position of the narrow line was taken as the rest frame reference for each transient.
@@
In the bottom left plot of Fig. 12, the flux of the broad H is shown. A progressive increase in the flux is observed in all SNe, with a peak between 200 and 300 after the explosion. This is followed by a decrease in the case of SNe 2021adxl and 2022wed, while it appears to remain constant (albeit within very large errorbars) for SN 2021acya. A high H luminosity is often correlated to strong interaction (Chugai, 1991). If this is true also for our SNe, the strength of interaction increases with time and peaks after the light curve peak. The epoch at which the H flux starts to decrease is likely related to the extent and density profile of the interacting CSM.
@@
Finally, the ratio of for both the narrow and the broad components is calculated. For the narrow line, the ratio is constant within the errors and close to the 3.1 limit value for the case B hydrogen recombination (Osterbrock & Ferland, 2006). For the narrow lines, this indicates that the level population is determined by the radiation field. The broad lines ratio is shown in the bottom right panel of Fig. 12, where we consider only ratios with an error below 25%. The values are significantly higher than for the narrow lines and the ratio changes over time. This happens when electrons are pushed to higher levels through collisions, in regions of high density and optical depth, confirming that all SNe in our sample have a strong shock component in their light emission (Chevalier & Fransson, 1994). This is also in agreement with what is found in the literature, for example, for SNe 2010jl and 2013L (Fransson et al., 2014; Taddia et al., 2020).
@@
Brennan et al. (2023) pointed to a possible two-component absorption feature in the He i P-Cygni profile of SN 2021adxl, which they explain as either a blend with the Na I D or the effect of an asymmetric explosion. In Fig. 13, a comparison of the spectrum of SN 2021adxl at +91 with the one of SN 2022wed at +136 is shown. The wavelength of the blue bump excess matches the one of SN 2022wed and the presence of Fe in this SN is confirmed by Fe II in multiple spectra. This shows that the possible high-velocity He i component mentioned by Brennan et al. (2023) is rather better understood as the emergence of the blue bump due to Fe ii emission (cf. Sect. 2.7). Instead, we agree on the interpretation of the lower velocity component as He i at . This feature will be discussed later on.
@@
When the shocked region is exposed, it is possible to measure the shock velocity from the blue shoulder of the H line. The CDS is confined between FS and RS and it is optically thick but radially thin. Photons emitted from such a structure will give a boxy profile to the line (Dessart et al., 2015) and the blueward limit of the emission is a direct measure of . Even in the case of a composite profile with electron scattering wings, it is still possible to recover the shock velocity as long as the blue shoulder is visible. As shock velocity, we took the Doppler shift of the blue shoulder compared to the position of the narrow emission. The measurements were performed on the spectra where the boxy component is visible and are shown in Fig. 14, while the highest and lowest values are reported in Table 4. The measurements have a significant scatter due to the uncertainty of the blue shoulder position, but it is still possible to fit them to a declining power-law, which gives . The same was done for SN 2022qml, finding , more similar to the velocity evolution Taddia et al. (2020) find for SN 2013L. The measured on the spectra are reported in Table 4, along with other parameters inferred from the light curves and spectra.
@@ 2.7 Blue bump excess
As we already noticed, a peculiar feature in the spectra of many interacting SNe is a blue bump bluewards from Å. This was first observed in SN 1988Z (Turatto et al., 1993) and it is typically attributed to a forest of Fe II lines (see, e.g., Foley et al. 2007; Smith et al. 2012a). The relative strength of this feature varies depending on the objects considered and the phases. In general, the spectrum of interacting SNe is a combination of a hot continuum and emission lines, some isolated and others, such as those contributing to the blue bump, blended. To estimate the relative contribution of the two components, we fit a BB function in the red part on the redshift-corrected spectra ( Å, onward), avoiding the most prominent emission lines (in particular, H). However, we caution that the BB fits fails after roughly 200 (100 for SN 2022qml). An example of the fit region and the estimated BB is shown in Fig. 17. Here, the black line is the observed spectrum of SN 2021acya at +61, while the red bands show the location of the fitting region and the green line the fitted BB function. A blue line representing the BB fit to the whole spectrum is also added to show that the fit is less optimal in this case due to the blue bump. The blue bump excess is defined as the ratio between the difference in flux between the observed flux in the blue (avoiding the most prominent emission lines, in particular, H) and the flux below the BB fitting on the redder part of the spectrum at the same wavelengths, all divided by the total BB flux.
@@
Considering the gaps in the observations and the different phases of the SNe in our sample, a mean blue bump excess was computed every 100 since the explosion to aid in the comparison. It is shown, compared to the mean luminosity at the same phases derived from the g,r,i pseudo-bolometric light curve (see Sect. 3 for details on the computation), in Fig. 18, where three reference objects are also added to represent the effect of different amounts of interaction: the strongly-interacting SN 2010jl, the mildly interacting SN 1998S, and the Type IIP SN 2004et, where moderate interaction only started after (Kotak et al., 2009). The evolution with time is significant. The excess is small (in some cases, even negative, meaning that the fitted BB exceeds the measured flux, probably due to line blanketing) at early phases, and tends to grow after roughly 100, while the luminosity decreases. SNe 2010jl and 2021adxl seem to show milder blue bump excesses at early phases than SNe 2021acya and 2022wed, while SN 2022qml is the most extreme, with a stark increase in the blue bump excess after 100. SN 1998S exhibits a similar behaviour, albeit with lower absolute values of the blue bump excess. SN 2004et, on the other hand, shows an always negative excess. This is in line with what we would expect from a non-interacting SN444Indeed SN 2004et underwent late rebrightening (phase ) due to interaction (Kotak et al., 2009). However, while almost all SNe are expected to undergo interaction, sooner or later, since all massive stars suffer from mass-loss episodes, the interaction happening at such late phases is orders of magnitude less strong than for strongly-interacting SNe, since the density of the CSM is lower..
@@ 3 Bolometric light curves
The SNe in our sample have different coverage in terms of wavelength and phases. To ensure a more objective comparison among them, we computed a pseudo-bolometric light curve for all SNe integrating the flux in the g,r,i filters. The magnitudes are corrected for extinction and converted to flux densities using photometric zero points555http://svo2.cab.inta-csic.es/theory/fps/. The flux is then integrated in the sampled spectral region using a trapezoidal rule and assuming zero flux below/above the limit defined by the filter equivalent width of the bluer/redder filter, respectively. The measured flux is converted into luminosity using the adopted distance modulus. The results for our four SNe is shown in Fig. 19, where the pseudo-bolometric light curves of SNe 2004et and 2010jl, calculated in the same way, are also included for comparison. A logarithmic scale is used to emphasise the change of slope between the early and late phases.
SN 2021acya is the most energetic SN in our sample, with an extremely high (Table 3). After the very bright peak, SN 2021acya shows a slow decline followed by a plateau (better seen in Fig. 1) that lasts from +150 days until +300 days. The decline rate both before and after the plateau is similar to that of SN 2010jl. SN 2021adxl shares some similarities with SN 2010jl, in particular, the break in the light curve around +400. However, the decline rate is much higher and, in fact, SN 2021adxl is about 1.5 mag fainter than SN 2010jl at . SN 2022qml has the shortest light curve, since it is observed for less than 200, after which the SN is lost because of conjunction with the Sun. When it re-emerged from behind the Sun, the SN was no longer detected. The decline rate is also similar to that of SN 2010jl. Finally, SN 2022wed has the most peculiar light curve shape, with its two peaks, the second much brighter than the first. It also does not show a break at 400, however, the decline rate after the second peak is almost identical to that of SN 2021acya after the plateau.
@@
Along with the direct comparison, we tried to gauge the true bolometric luminosity taking into account UV and NIR bands. In this respect, SN 2021acya has the best photometric coverage (see Table 6). To estimate the contribution of UV and NIR photometry, in Fig. 20 the bolometric light curve computed with the contribution from the UV or NIR flux is compared to the one computed using only g,r,i. The contributions are calculated as the ratio between the UV or NIR flux and the total flux. While the NIR contribution at early phases is small (around 20%) and rises up to 40% at later times, the UV contribution is significant, up to 60%, but then it decreases rapidly. A similar UV excess was also measured in SN 2010jl and other interacting SNe (Fransson et al., 2014). In SN 2010jl, the early NIR light curves followed the decline of the optical ones, while after they flattened in and and even increased the flux in . This was attributed to pre-existing dust in the CSM that was re-heated by the optical emission (Fransson et al., 2014). For SN 2010jl, the NIR bands contribute 20 to 50% of the bolometric flux in the phase range . This is in line with the NIR contribution that we measure after the maximum for SN 2021acya.
The same check was also performed on SNe 2021adxl (for which we have a sequence of u,z,J,H,K observations, see Table 7), 2022qml (for which we have a sequence of observations, see Table 8), and 2022wed ( observations, see Table 9). The band contribution is consistent in all SNe, adding alone between % of the total flux at early phases. The NIR observations of SN 2021adxl, on the other hand, behave like in SN 2021acya, giving a flux % higher. Also, the NIR contribution appears to increase with time.
@@
In Khatami & Kasen (2023), the authors explore the variability of SN light curves when interaction with the CSM is involved. The key parameter in their model is the break-out parameter :
@@ 4 Mass-loss rate and mass of the CSM
Even with all the caveats due to the assumptions described above, the values of and are much higher than those expected for a steady wind and more in line with eruptive processes (Matsumoto & Metzger, 2022), in agreement with an LBV progenitor (for reference, Fassia et al. 2001 find for SN 1998S a mass-loss rate of , two orders of magnitude lower than our values, which are instead more in line with the eruptive episodes in Eta Carinae). To produce a CSM mass of requires a huge progenitor, the formation of which could be challenging for stellar evolution models. The only SN in our sample with a considerably smaller is SN 2022qml, once again fortifying the evidence of a different progenitor, perhaps, even a thermonuclear explosion embedded in a H-rich CSM.
It is also interesting to explore the mass-loss rate as a function of time. In particular, assuming a steady wind, it is possible to derive the value of on the timescale of the wind. The results for each SN are shown in Fig. 21. Here, there is a strong dependence especially on the ratio between the shock and the wind velocity, which gives the overall shape to the trend, as well as the duration of the bolometric light curve. As shock velocity, we considered the trend derived on the H blue shoulder fit for SNe 2021adxl and 2022qml, while for SNe 2021acya and 2022wed we take the same trend of SN 2021adxl rescaled to the velocity derived by Fransson et al. (2014) at 320 days for SN 2010jl, as explained in Sect. 2.6. From Fig. 21, it appears that the mass-loss rate of SN 2022qml was more constant, compared to the other SNe, and overall lower, once again in agreement with a different origin for this SN. The trends of SNe 2021adxl and 2022wed are similar but shifted, with SN 2022wed starting to significantly increase later on, around 20 years after the onset, while SN 2021adxl shows a high mass-loss rate already 10 years in. SN 2021acya, on the other hand, has a peculiar behaviour with possibly three distinct peaks that could indicate the onset of different mass-loss episodes. Interestingly, the timeline of the peaks seems to align with the other SNe: the second one happens around the same time as the one of SN 2021adxl and the third is almost in coincidence with the peak of SN 2022wed. With these considerations, it appears even more evident that the progenitor stars of these SNe had to be massive, peculiar objects to be able to expel such a huge amount of mass within such a limited time period, possibly, in the case of SN 2021acya, in two separate episodes. A mass-loss rate that is high for decades and increases with the approaching of the explosion is also in line with what is found by Moriya et al. (2014).
@@
For reference, we can apply the same analysis also to SN 2020faa, which has a derived kinetic energy and CSM mass and radius: , , and (Salmaso et al., 2023). The small CSM radius disfavours SN 2020faa as a likely HE neutrino source, however, the kinetic energy and the CSM mass are in agreement with what is found for the other SNe in our sample. As mentioned, degeneracy in the parameters implies uncertainties that are difficult to disentangle but it is clear that high values of CSM mass and kinetic energy, hence requiring a massive progenitor, seem to be needed.

### v2

@@ 2.6 Analysis on the line profile and emission
During the interaction between the SN ejecta and the CSM, four main regions should be considered: from the outside, i) the unshocked CSM, ii) the CSM that was shocked by a forward shock (FS), iii) the SN ejecta shocked by the reverse shock (RS), and iv) the unshocked SN ejecta that are expanding fast (Chevalier & Fransson, 1994). If the CSM is dense (), a cool dense shell (CDS) forms between the FS and RS (Chevalier & Fransson, 1985). An opaque CDS has the same effect of an expanding photosphere (Chugai, 2001). In the case of strong interaction, the heated CSM dominates the optical emission. Since emissions from the CSM can remain bright for months or years, it is possible that the ejecta thermal energy fades before the CSM becomes transparent and thus one may never see the characteristic P-Cygni profiles indicative of an expanding photosphere (Smith, 2017). In general, in SNe IIn one may expect a narrow () component due to the unshocked CSM and a broader () component which, depending on the optical depth of the CSM can originate either from the ionised SN ejecta or from electron-scattering in the CSM (Huang & Chevalier, 2018). Furthermore, if the CSM is asymmetric (Stritzinger et al., 2012), or if dust is present (Fox et al., 2011), they can affect the shape of the line profiles, thus making the theoretical interpretation more difficult.
@@
In Fig. 11, we plot a zoom on the H region of spectra taken around 100 from the explosion for all the SNe in our sample. SNe 2021adxl and 2022qml have broad, blue, asymmetric profiles with a narrow emission centered at rest wavelength. On the other hand, SNe 2021acya and 2022wed have a more symmetric broad line (Tables and are available online only, see the Data availability section). In the case of SN 2021acya, there is a small dip that may be a narrow P-Cygni absorption but it is probably an instrumental artifact, since it is not visible at other epochs.
@@
Given the diversity in shape of the H across our sample, it is difficult to directly compare them in terms of FWHM, line position, and total flux. Opting to apply the same method to all of them while being conscious that the procedure will not be optimal, we performed a multi-component fit for the H lines, including the contribution of a broad Gaussian and a narrow Lorentzian function. We chose these rather simple profiles because they can ensure a more direct comparison among all the SNe in the sample, even if they might not provide the most perfect fit, especially for complicated profiles such as those of SNe 2021adxl and 2022qml. In some spectra, adding a second, intermediate-width Gaussian component representing the post-shock gas (Smith, 2017) allowed for a better fit. This is a relatively faint feature that was ignored for the rest of the analysis. For each component, we fit the position, peak intensity, and width. Given the asymmetric profile of the line in some objects, the position of the peak of the broad component was allowed to vary with respect to the narrow one. All the measurements were performed on reddening-corrected spectra.
@@
For all the SNe in the sample, position, FWHM, and intensity of the narrow component are constant within the errors (calculated by summing in quadrature the root mean square and the resolution). Also, the FWHM of the narrow component is not resolved in the fit with the exception of the spectrum of SN 2022qml at +49, which yields a FWHM of 12.9 Å. The average position of the narrow line was taken as the rest frame reference for each transient.
@@
In the bottom left plot of Fig. 12, the flux of the broad H is shown. A progressive increase in the flux is observed in all SNe, with a peak between 200 and 300 after the explosion. This is followed by a decrease in the case of SNe 2021adxl and 2022wed, while it appears to remain constant (albeit within very large errorbars) for SN 2021acya. A high H luminosity is often correlated to strong interaction (Chugai, 1991). If this is true also for our SNe, the strength of interaction increases with time and peaks after the light curve peak. The epoch at which the H flux starts to decrease is likely related to the extent and density profile of the interacting CSM. A more extended CSM would sustain the interaction for a longer time, while a rapidly declining density would drop the strength of interaction and thus the luminosity. In this respect, the decrease for SN 2022wed happens earlier than for the others, around the maximum of the second peak. Possibly, this denotes a rapid change in the CSM density for this particular object. On the other hand, for SNe 2021acya and 2021adxl the change happens later on, perhaps indicating a less steep decline in the CSM density. Also, note that the total H luminosity is higher for SN 2021acya compared to SN 2021adxl, possibly, due to a higher CSM mass for the former.
@@
Finally, the ratio of for both the narrow and the broad components is calculated. For the narrow line, the ratio is constant within the errors and close to the 3.1 limit value for the case B hydrogen recombination (Osterbrock & Ferland, 2006). For the narrow lines, this indicates that H is excited by radiation. The broad lines ratio is shown in the bottom right panel of Fig. 12, where we consider only ratios with an error below 25%. The values are significantly higher than for the narrow lines and the ratio changes over time. This happens when electrons are pushed to higher energy levels through collisions, in regions of high density and optical depth, confirming that all SNe in our sample have a strong shock component in their light emission (Chevalier & Fransson, 1994). This is also in agreement with what is found in the literature, for example, for SNe 2010jl and 2013L (Fransson et al., 2014; Taddia et al., 2020).
@@
Brennan et al. (2024) pointed to a possible two-component absorption feature in the He i P-Cygni profile of SN 2021adxl, which they explain as either a blend with the Na I D or the effect of an asymmetric explosion. In Fig. 13, a comparison of the spectrum of SN 2021adxl at +91 with the one of SN 2022wed at +136 is shown. The starting position of the blue bump excess matches the one of SN 2022wed and the presence of Fe in this SN is confirmed by Fe II in multiple spectra. This shows that the possible high-velocity He i component mentioned by Brennan et al. (2024) is rather better understood as the emergence of the blue bump due to Fe ii emission (cf. Sect. 2.7). Instead, we agree on the interpretation of the lower velocity component as He i at . This feature will be discussed later on.
@@
When the shocked region is exposed, it is possible to measure the shock velocity from the blue shoulder of the H line. The CDS is confined between FS and RS and it is optically thick but radially thin. Photons emitted from such a structure will give a boxy profile to the line (Dessart et al., 2015) and the blueward limit of the emission is a direct measure of . Even in the case of a composite profile with electron scattering wings, it is still possible to recover the shock velocity as long as the blue shoulder is visible. As shock velocity, we took the Doppler shift measured on the left corner of the blue shoulder compared to the position of the narrow emission. The measurements were performed on the spectra where the boxy component is visible and are shown in Fig. 14, while the highest and lowest values are reported in Table 4. The measurements have a significant scatter due to the uncertainty of the blue shoulder position, but it is still possible to fit them to a declining power-law, which gives . The same was done for SN 2022qml, finding , more similar to the velocity evolution Taddia et al. (2020) find for SN 2013L. The measured on the spectra are reported in Table 4, along with other parameters inferred from the light curves and spectra.
@@ 2.7 Blue bump excess
As we already noticed, a peculiar feature in the spectra of many interacting SNe is a blue bump bluewards from Å. This was first observed in SN 1988Z (Turatto et al., 1993) and it is typically attributed to a forest of Fe II lines (see, e.g., Foley et al. 2007; Smith et al. 2012a). The relative strength of this feature varies depending on the objects considered and the phases. In general, the spectrum of interacting SNe is a combination of a hot continuum and emission lines, some isolated and others, such as those contributing to the blue bump, blended. To estimate the relative contribution of the two components, we fit a black body (BB) function in the red part on the redshift-corrected spectra ( Å, onward), avoiding the most prominent emission lines (in particular, H). However, we caution that the BB fits fails after roughly 200 (100 for SN 2022qml). An example of the fit region and the estimated BB is shown in Fig. 17. Here, the black line is the observed spectrum of SN 2021acya at +61, while the red bands show the location of the fitting region and the green line the fitted BB function. A blue line representing the BB fit to the whole spectrum is also added to show that the fit is less optimal in this case due to the blue bump. The blue bump excess is defined as the ratio between the difference in flux between the observed flux in the blue (avoiding the most prominent emission lines, in particular, H) and the flux below the BB fitting on the redder part of the spectrum at the same wavelengths, all divided by the total BB flux.
@@
Considering the gaps in the observations and the different phases of the SNe in our sample, a mean blue bump excess was computed every 100 since the explosion to aid in the comparison. It is shown, compared to the mean luminosity at the same phases derived from the g,r,i pseudo-bolometric light curve (see Sect. 3 for details on the computation), in Fig. 18, where three reference objects are also added to represent the effect of different amounts of interaction: the strongly-interacting SN 2010jl, the mildly interacting SN 1998S, and the Type IIP SN 2004et, where moderate interaction only started after (Kotak et al., 2009). The evolution with time is significant. The excess is small (in some cases, even negative, meaning that the fitted BB exceeds the measured flux, probably due to line blanketing) at early phases, and tends to grow after roughly 100, while the luminosity decreases. SNe 2010jl and 2021adxl seem to show milder blue bump excesses at early phases than SNe 2021acya and 2022wed, while SN 2022qml is the most extreme, with a stark increase in the blue bump excess after 100. SN 1998S exhibits a similar behaviour, albeit with lower absolute values of the blue bump excess. SN 2004et, on the other hand, shows an always negative excess. This is in line with what we would expect from a non-interacting SN666Indeed SN 2004et underwent late rebrightening (phase ) due to interaction (Kotak et al., 2009). However, while almost all SNe are expected to undergo interaction, sooner or later, since all massive stars suffer from mass-loss episodes, the interaction happening at such late phases is orders of magnitude less strong than for strongly-interacting SNe, since the density of the CSM is lower..
@@ 3 Bolometric light curves
The SNe in our sample have different coverage in terms of wavelength and phases. To ensure a more objective comparison among them, we computed a pseudo-bolometric light curve for all SNe integrating the flux in the g,r,i filters. The magnitudes are corrected for extinction and converted to flux densities using photometric zero points777http://svo2.cab.inta-csic.es/theory/fps/. The flux is then integrated in the sampled spectral region using a trapezoidal rule and assuming zero flux below/above the limit defined by the filter equivalent width of the bluer/redder filter, respectively. The measured flux is converted into luminosity using the adopted distance modulus. The results for our four SNe is shown in Fig. 19, where the pseudo-bolometric light curves of SNe 2004et and 2010jl, calculated in the same way, are also included for comparison. A logarithmic scale is used to emphasise the change of slope between the early and late phases.
SN 2021acya is the most energetic SN in our sample, with an extremely high (Table 3). After the very bright peak, SN 2021acya shows a slow decline followed by a plateau (better seen in Fig. 1) that lasts from +150 days until +300 days. The decline rate both before and after the plateau is similar to that of SN 2010jl. SN 2021adxl shares some similarities with SN 2010jl, in particular, the break in the light curve around +400. However, the decline rate is much higher and, in fact, SN 2021adxl is about 1.5 mag fainter than SN 2010jl at . SN 2022qml has the shortest light curve, since it is observed for less than 200, after which the SN is lost because of conjunction with the Sun. When it re-emerged from behind the Sun, the SN was no longer detected. The decline rate is also similar to that of SN 2010jl. Finally, SN 2022wed has the most peculiar light curve shape, with its two peaks, the second much brighter than the first. This is probably due to a CSM arranged in two shells with different mass and density. It also does not show a break at 400, however, the decline rate after the second peak is almost identical to that of SN 2021acya after the plateau.
@@
Along with the direct comparison, we tried to gauge the true bolometric luminosity taking into account UV and NIR bands. In this respect, SN 2021acya has the best photometric coverage. To estimate the contribution of UV and NIR photometry, in Fig. 20 the bolometric light curve computed with the contribution from the UV or NIR flux is compared to the one computed using only g,r,i. The contributions are calculated as the ratio between the UV or NIR flux and the total flux. While the NIR contribution at early phases is small (around 20%) and rises up to 40% at later times, the UV contribution is significant, up to 60%, but then it decreases rapidly. A similar UV excess was also measured in SN 2010jl and other interacting SNe (Fransson et al., 2014). In SN 2010jl, the early NIR light curves followed the decline of the optical ones, while after they flattened in and and even increased the flux in . This was attributed to pre-existing dust in the CSM that was re-heated by the optical emission (Fransson et al., 2014). For SN 2010jl, the NIR bands contribute 20 to 50% of the bolometric flux in the phase range . This is in line with the NIR contribution that we measure after the maximum for SN 2021acya.
The same check was also performed on SNe 2021adxl (for which we have a sequence of u,z,J,H,K observations), 2022qml (for which we have a sequence of observations), and 2022wed ( observations). The band contribution is consistent in all SNe, adding alone between % of the total flux at early phases. The NIR observations of SN 2021adxl, on the other hand, behave like in SN 2021acya, giving a flux % higher. Also, the NIR contribution appears to increase with time.
@@
In Khatami & Kasen (2023), the authors explore the diversity of SN light curves when interaction with the CSM is involved. The key parameter in their model is the break-out parameter :
@@ 4 Mass-loss rate and mass of the CSM
Even with all the caveats due to the assumptions described above, the values of and are much higher than those expected for a steady wind and more in line with eruptive processes (Matsumoto & Metzger, 2022), in agreement with an LBV progenitor (for reference, Fassia et al. 2001 find for SN 1998S a mass-loss rate of , two orders of magnitude lower than our values, which are instead more in line with the eruptive episodes in Eta Carinae). To produce a CSM mass of requires a huge progenitor, the formation of which could be challenging for stellar evolution models. The only SN in our sample with a considerably smaller is SN 2022qml. This, together with the different spectra and shorter light curve duration, points toward a possible thermonuclear explosion whose progenitor was embedded in a H-rich CSM.
It is also interesting to explore the mass-loss rate as a function of time. Regardless of the CSM origin (wind or eruption), we assume that it is distributed in a shell within which the velocity is constant. Assuming a steady wind for simplicity, it is possible to derive the value of on the timescale of the wind. The results for each SN are shown in Fig. 21. Here, there is a strong dependence especially on the ratio between the shock and the wind velocity, which gives the overall shape to the trend, as well as the duration of the bolometric light curve. As shock velocity, we considered the trend derived on the H blue shoulder fit for SNe 2021adxl and 2022qml, while for SNe 2021acya and 2022wed we take the same trend of SN 2021adxl rescaled to the velocity derived by Fransson et al. (2014) at 320 days for SN 2010jl, as explained in Sect. 2.6. From Fig. 21, it appears that the mass-loss rate of SN 2022qml was more constant, compared to the other SNe, and overall lower, once again in agreement with a different origin for this SN. The trends of SNe 2021adxl and 2022wed are similar but shifted, with SN 2022wed starting to significantly increase later on, around 20 years after the onset, while SN 2021adxl shows a high mass-loss rate already 10 years in. SN 2021acya, on the other hand, has a peculiar behaviour with possibly three distinct peaks that could indicate the onset of different mass-loss episodes. Interestingly, the timeline of the peaks seems to align with the other SNe: the second one happens around the same time as the one of SN 2021adxl and the third is almost in coincidence with the peak of SN 2022wed. With these considerations, it appears even more evident that the progenitor stars of these SNe had to be massive, peculiar objects to be able to expel such a huge amount of mass within such a limited time period, possibly, in the case of SN 2021acya, in two separate episodes. A mass-loss rate that starts high and increases for decades after the onset of the episode is also in line with what is found by Moriya et al. (2014), although it may decline closer to the explosion.
@@
For reference, we can apply the same analysis also to SN 2020faa, a SN powered by hidden interaction, which has a derived kinetic energy and CSM mass and radius: , , and (Salmaso et al., 2023). The small CSM radius disfavours SN 2020faa as a likely HE neutrino source, however, the kinetic energy and the CSM mass are in agreement with what is found for the other SNe in our sample. As mentioned, degeneracy in the parameters implies uncertainties that are difficult to disentangle but it is clear that high values of CSM mass and kinetic energy, hence requiring a massive progenitor, seem to be needed.